OpenStudy (anonymous):

Medal for help - what is the value of x/y when x =9/4 and y=3/5? a. 15/4 b. -15/4 c.4/3 d. 27/20

OpenStudy (anonymous):

@AnswerMyQuestions @greenlegodude57

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

b

OpenStudy (anonymous):

@d3Xter

OpenStudy (anonymous):

@davyjones

OpenStudy (kirbykirby):

\[\large \frac{\frac{9}{4}}{\frac{3}{5}}=\frac{9}{4}\div\frac{3}{5}=\frac{9}{4}\times\frac{5}{3}\]

OpenStudy (anonymous):

a

OpenStudy (anonymous):

A ? does every agree ?

OpenStudy (kirbykirby):

yes A

OpenStudy (anonymous):

@kirbykirby can u give davyjones a medal too please ty

OpenStudy (kirbykirby):

:)